5 mL of nitric acid, HNO3, are required to react with 50 mL of a 1.25 M Barium hydroxide, Ba(OH)2 solution:


2 HNO3 (aq) + Ba(OH)2 (aq) --> 2 H2O (aq) + Ba(NO3)2 (aq)



Determine the concentration (Molarity) of the nitric acid solution (HNO3).

1
Expert's answer
2022-02-17T09:50:03-0500

1ml=1g

Number of moles of nitric acid;

Molar mass= 63

Moles= 5/63= 0.0794moles

Molarity= moles/ litre


Molarity= 0.0794/1000

= 7.9× 10^-5M